OpenStudy (anonymous):

Help,. how to find the cube root of (1.728)(10^-6)???

OpenStudy (turingtest):

\[\sqrt[3]{10^{-6}}=10^{-6/3}\]^^^ found through algebra\[\sqrt[3]{1.728}\]^^^ found through calculator.

OpenStudy (phi):

isn't it common knowledge that 12^3= 1728?

OpenStudy (turingtest):

even better, thanks phi. need to review my cubes...

OpenStudy (anonymous):

so the answer is (1.2)(10^-2) ? can it be in the form: \[(1.2)/(10^{2})\] ?

OpenStudy (phi):

write 1.728(10^-6) as 1728 x 10^-3 x 10^-6 = 1728x10^-9 cube root is 12x10^-3= 1.2(10^-2) or 0.012

OpenStudy (phi):

which is the same as your answer

OpenStudy (anonymous):

aw... thanks phi!